OpenStudy (anonymous):

How many terms of the series do we need to add in order to find the sum to the indicated accuracy? Note: Enter the smallest possible integer. sum n=1 to infinity (-1)^n-1(7/n^4), error <= .001 (This is an alternating series).

OpenStudy (blockcolder):

\[\sum_{n=1}^{\infty}(-1)^{n-1}\frac{7}{n^4} \qquad b_n=\frac{7}{n^4}\] The error of an alternating series at the nth partial sum is less than the (n+1)th term. Thus, to ensure that the error is less than 0.001, we should have \[0.001\leq \frac{7}{(n+1)^4}\] Solve for n here then round up the value to ensure the desired accuracy. :D